91 ** A U-tube contains liquid of unknown specific grav- ity. An oil of density 800 kg/m3 is poured into one arm of the tube until the oil column is 12 cm high. The oil-air interface is then 5.0 cm above the liquid level in the other arm of the U-tube. Find the specific gravity of the liquid. when empty is 32 kg. The crate contains gold bullion that fills 36 percent of its volume; the remaining volume is filled with seawater. (a) What is the tension in the cable that raises the crate and bullion while the crate is below the surface of the sea? (b) What is the tension in the cable while the crate is lifted to the deck of the ship if (1) none of the seawater leaks out of the crate, and (2) the crate is lifted so slowly that all of the sea- water leaks out of the crate?
Answer #1

Using the equation,

p = p0 + ρgh

pi = pressure at interface of the two liquids.

pA = atmospheric pressure. ρ1 = density of oil. ρ2 = density of unknown liquid. h1 = 12cm. h2 = 5cm.

At the interface between the oil and liquid the pressure will be the same as the pressure 12 - 5 = 7cm below the surface:

pA + ρ2g(h1 - h2) = pA + ρ1gh1

ρ2 = ρ1h1 / (h1 - h2) = [800kg/m^3][12/7] = 1371.4285 kg/m^3

Second question is incomplete, as per HOMEWORKLIB POLICY I have solved one complete question.